# Examining the Prevalence and Dynamics of AI-Generated Media in Art Subreddits

### v2

@@ 1. Introduction
We are interested in how generative AI changes some core aspects of online community life. The first aspect is content production. Online communities lean on the active production of relevant content through posts and discussions to fulfill their purpose and maintain their perceived popularity (Gero et al., 2023; Panchanadikar et al., 2024; Atcheson et al., 2024). With generative AI, we may expect shifts in content production, both in terms of quantity and perceived quality, as AI tools change who can create and what they create (Wei and Tyson, 2024; Zhou and Lee, 2024). The second aspect is participation. Online communities rely on participation, and in this regard, practices, policies, and norms around participation barriers often dictate the dynamics and even success of communities (Atcheson et al., 2024; Frosio, 2023; Radivojevic et al., 2024; Lloyd et al., 2025a). AIGC may influence who participates, potentially lowering barriers for newcomers or conversely, deterring some contributors (Kraut et al., 2012; Wei and Tyson, 2024). The third aspect is the communication and trust between community members. In online communities, it is often critical that members trust each other and have a positive exchange (Mazzone and Elgammal, 2019). AIGC may alter how members engage with each other and the overall social dynamics in the community (Guo et al., 2024; Garcia, 2024). We therefore address the following research questions:
@@ 2.3. AI-generated Art and Online Communities
The papers mentioned above examined dedicated art communities, with centralized policies and rules. In this work, we are focused on Reddit, and particularly on the decentralized, mostly-self governed (Fiesler et al., 2018) communities on this site, and especially those dedicated to art. Recent work has shown that subreddit moderators are already grappling with AIGC (Lloyd et al., 2025b, a). Subreddit moderators are creating rules that regulate AI use in their communities, clarify norms, and enforce policies. By leveraging the organizational structure of Reddit’s subreddits, we are able to analyze data based on the distinct characteristics of different communities, including their own rules — namely, if the community explicitly disallows AI, or is neutral towards it in its rules. Our analysis also covers questions such as submissions volume and number of participants, in light of the different community types.
@@ 4. Classification of the Dataset Using LLMs
Having gathered a dataset of potentially AI-relevant image-based posts based on keyword match, we further classified the posts using large language models (LLMs) (Abburi et al., 2023; Zhang et al., 2024; Pangakis and Wolken, 2024). First, since keyword matching can be imprecise, we use classifiers to filter the dataset to only retain posts and comments genuinely pertaining to AI-generated images. Second, we differentiate between text that identifies the image in the post as being AI-generated, versus those that broadly mention AI-generated images.
@@
To this end, we developed textual prompts and utilized the ChatGPT API [^5] https://platform.openai.com/docs/api-reference/introduction (GPT-4o model). Three distinct prompts were developed to correspond with each of the three type of text in our dataset: post’s title or body, comments made by the author of the post, and comments by other users (see Appendix C). These prompts were designed following the principles of chain-of-thought prompting (Wei et al., 2022). Each prompt required the language model to provide reasoning before giving a classification decision. Additionally, each prompt included nine examples from the dataset, implementing an in-context few-shot learning approach (Wei et al., 2022).
@@
We evaluated the accuracy of each of the three prompts using three validation sets. We randomly sampled 100 texts from each type, and manually labeled them according to the established coding. That is, the texts were labeled as follows: (A) Implies AI involvement, when the text explicitly or implicitly suggests that AI tools were involved in generating the attached image; (B) Does not imply AI, when the text mentions AI-generated images, but does not imply that the attached image is AI-generated (including texts that suggests that the attached image is not AI-generated); or (C) Unrelated to AI, only when the text is completely unrelated to AI-generated images (regardless of whether the text refers to the image in the post)—in most cases, a false positive of the keyword match. Next, we classified the three validation sets using the corresponding prompts. The prompts, along with examples of input texts and their corresponding classifications, are provided in Appendix C. This classification task achieved an accuracy of 93% for the title of the posts, 82% for comments made by the author of the post, and 84% for comments from other accounts. In comparison, the baseline accuracy for the majority class is 87% for post titles, 49% for comments made by the post’s author, and 55% for comments from other accounts. Note that the majority class differs across the three validation sets: for post titles and comments made by the post’s author, the majority label is Implies AI involvement, while for comments from other accounts, the majority label is Does not imply AI.
@@
We classified all data entities in the dataset using these prompts, following a process inspired by the QuaLLM framework (Rao et al., 2024). According to this process we submitted the prompt, consisting in each case of the relevant class explanation and the in-context examples, with every 100 input texts at a time [^6] Note that during the evaluation phase we also compared the accuracy values obtained by submitting different numbers of entities as input (i.e., 10 instead of 100) and achieved identical results.. Similarly, we requested that the LLM return the final output as a dictionary, including both the reasoning and classifications for all these 100 texts.
@@ 6.2. Participation and Newcomers
The previous sections focused on the volume of posts and discussions over time and in relation to other image posts in the same communities. In this section, we address our second research question by examining participation patterns in the creation of AI-related posts across the different art communities. In particular, we focus on newcomers (first-time image posters). Is there evidence that AI-generated images make it easier for newcomers to contribute content – and if so, do these newcomers stay and contribute over time? Does this pattern diverge for those who are making Transparent AI posts to AI-disallowed versus AI-neutral communities?
Overall, the participation of newcomers was significantly more emphasized in making Transparent AI posts than the overall presence of newcomers in posting images to these communities. In AI-disallowed communities, newcomers usually publish 6.71% of all image-based posts. This ratio is similar in AI-neutral communities, with 6.51% of newcomers publishing image-based posts. However, newcomers made as much as 25.15% of Transparent AI posts in AI-disallowed communities during our period of analysis. In AI-neutral communities, newcomers made a similar 22.99% of the Transparent AI posts. The differences between the newcomers’ rate of production of identified AI content versus regular content was significant for both types of communities (Chi-square, $p<0.00001$). This over-representation in newcomers posting AI-generated content suggests that this kind of content may encourage and broaden community participation.
@@
Overall, the participation trends shows that newcomers more heavily rely on posting AI content to participate in these communities, but that (as expected) their participation does not last long in AI-disallowed communities. These trends may relate to the type of interactions around AI-generated content, which we look at next.
@@ 6.3. Feedback, Interactions and Sentiment
The time dynamics show a different response to suspected images in the two types of communities. To understand the response better, we conducted a sentiment analysis of comments suggesting AI involvement in the Suspected AI posts. To label the sentiment of comments implying AI, we used a large language model as a classifier, following a similar approach to that described in the previous sections, which is based on prior studies (Wei et al., 2022; Rao et al., 2024). We used the LLM classifier to label the comments as Negative, Neutral, or Positive, using a prompt that can be found in Appendix D, along with examples of input texts and their corresponding classifications. To assess the accuracy of this method, we manually labeled a random sample of 100 comments and achieved an accuracy rate of 92% for this sentiment classification task.
Using these labels, we looked at differences in comment sentiment between AI-disallow and AI-neutral communities, and how they may have changed over time. To do that, we compared sentiment in two distinct time periods: January-March 2023 and October-December 2023 (the end of our data). In January-March 2023, the percentage of negative sentiment in comments in AI-disallowing communities was 39% (6% were positive, and the rest neutral). The ratio of negative sentiment was lower in AI-neutral communities, at around 28%. By October-December 2023, however, the pattern reversed, with AI-neutral communities exhibiting a higher percentage of negative comments (51%) than AI-disallowing communities (42%). The reason for the flip could be the better control exercised by the AI-disallowing communities about AI posts. Still, the data shows an increase over time in the proportion of comments expressing negative sentiment among those claiming AI involvement, across both community categories.
@@ 7. Discussion
The data processing and analysis steps may also have certain limitations, which could be addressed in future research. Identifying posts containing AI-generated media relies on text-based methods, which sometimes may capture user opinion and reactions, rather than accurately reflect the actual prevalence of AI-generated content. For example, in AI-neutral subreddits where AI-generated content is permitted, users can be less inclined to explicitly mention it in posts or comments. In contrast, in AI-disallowed subreddits, users may be more likely to claim that others’ content is AI-generated as it violates community rules. In the categorization of the communities step, given that AI-neutral communities do not explicitly address AI usage in their guidelines or descriptions, while AI-disallowed communities explicitly prohibit it, the observed differences in trends between these groups may reflect broader factors, such as more adaptive approach to emerging technologies by the AI-disallowed communities. That is, factors beyond the subreddit-specific policies regarding AI-generated content. Additionally, the analysis may face potential limitations related to the identification of newcomers, as the dataset may include bot or troll accounts considered as newcomers. Bots, in particular, are more likely to use AI-generated content (Grimme et al., 2022; Ferrara, 2023), and may be at higher risk of suspension by the platform. During data collection, measures were taken to filter out potential bot accounts based on usernames and repeated comments. However, these efforts may not have entirely excluded all accounts exhibiting bot-like behavior.

### v3

@@ 1. Introduction
We are interested in how generative AI changes some core aspects of online community life. The first aspect is content production. Online communities lean on the active production of relevant content through posts and discussions to fulfill their purpose and maintain their perceived popularity (Gero et al., 2023; Panchanadikar et al., 2024; Atcheson et al., 2024). With generative AI, we may expect shifts in content production, both in terms of quantity and perceived quality, as AI tools change who can create and what they create (Wei and Tyson, 2024; Zhou and Lee, 2024). Establishing the prevalence of AI-generated posts in online communities is crucial to understand the broader impact of these technologies in online social experiences. The second aspect is participation. Online communities rely on participation, and in this regard, practices, policies, and norms around participation barriers often dictate the dynamics and even success of communities (Atcheson et al., 2024; Frosio, 2023; Radivojevic et al., 2024; Lloyd et al., 2025a). AIGC may influence who participates, potentially lowering barriers for newcomers or conversely, deterring some contributors (Kraut et al., 2012; Wei and Tyson, 2024). The third aspect is the communication and trust between community members. In online communities, it is often critical that members trust each other and have a positive exchange (Mazzone and Elgammal, 2019). AIGC may alter how members engage with each other and the overall social dynamics in the community (Guo et al., 2024; Garcia, 2024). We therefore address the following research questions:
@@ 2.3. AI-generated Art and Online Communities
The papers mentioned above examined stand-alone online communities dedicated to art, with centralized policies and rules. In this work, we focus on Reddit, and particularly on the multiple decentralized, mostly-self governed (Fiesler et al., 2018) communities on this site, and especially those dedicated to art. Recent work has shown that subreddit moderators are already grappling with AIGC (Lloyd et al., 2025b, a). Subreddit moderators are creating rules that regulate AI use in their communities, clarify norms, and enforce policies. By leveraging the organizational structure of Reddit’s subreddits, we can analyze data based on the distinct characteristics of different communities, including their own rules — namely, if the community explicitly disallows AI, or is neutral towards it in its rules. Our analysis also covers questions around how AIGC may impact submission volume, the type of participants that contribute to online communities, and the community response to AIGC content in relation to different community moderation strategies.
Overall, while prior work emphasized moderation approaches and governance mechanisms (Lloyd et al., 2025b, a), our work complements these approaches by examining how AIGC interacts with the experience and practices of users – specifically how users disclose, perceive, and engage with such content.
@@ 4. Classification of the Dataset Using LLMs
Having gathered a dataset of potentially AI-relevant image-based posts based on keyword match, we further classified the posts using large language models (LLMs) (Abburi et al., 2023; Zhang et al., 2024; Pangakis and Wolken, 2024). The overall goal of this classification task is to identify (a) Transparent AI posts, where the creator implies the image is AI-generated (b) Suspected AI posts, where a commenter implies the image is AI-generated. We approach this task from two aspects. First, since keyword matching can be imprecise, we use classifiers to filter the dataset to only retain posts and comments genuinely pertaining to AI-generated images. Second, we differentiate between text that identifies the image in the post as being AI-generated, versus those that broadly mention AI-generated images. Importantly, we do not attempt to establish whether the image is actually AI-generated, since this task is beyond the scope of this paper which focuses on disclosures of and responses to AI-alleged images. While a perfect classifier for AI-generated images and art (a famously difficult challenge (Ghiurău and Popescu, 2024; Epstein et al., 2023)) would have been useful to substantiate these author and commenter claims about AI, the existence of such classifier would have negated the need to study disclosure and accusations altogether.
@@
Prior work has demonstrated the suitability of LLMs for text classification, when validated by human annotators (Choksi et al., 2024; Dunivin, 2025). We thus developed textual prompts and utilized the ChatGPT API [^5] https://platform.openai.com/docs/api-reference/introduction (GPT-4o model). Three distinct prompts were developed to correspond with each of the three types of text in our dataset: post’s title or body, comments made by the author of the post, and comments by other users (see Appendix C). These prompts were designed following the principles of chain-of-thought prompting (Wei et al., 2022). Each prompt required the language model to provide reasoning before giving a classification decision. Additionally, each prompt included nine examples from the dataset, implementing an in-context few-shot learning approach (Wei et al., 2022).
@@
We used the LLMs to label each text according to one of three categories: (A) Implies AI involvement, when the text explicitly or implicitly suggests that AI tools were involved in generating the attached image; (B) Does not imply AI, when the text mentions AI-generated images, but does not imply that the attached image is AI-generated (including texts that suggests that the attached image is not AI-generated); or (C) Unrelated to AI, only when the text is completely unrelated to AI-generated images (regardless of whether the text refers to the image in the post)—in most cases, a false positive of the keyword match. The prompts, along with examples of input texts and their corresponding classifications, are provided in Appendix C.
Following prior work (Choksi et al., 2024), we manually evaluated the accuracy of each of the three prompts using three validation datasets. We randomly sampled 100 texts from each of the three types of text, and manually labeled the resulting 300 texts according to the established coding. Next, we classified the three validation datasets using the corresponding prompts. The classifiers achieved an accuracy of 93% for the title of the posts, 82% for comments made by the author of the post, and 84% for comments from other accounts. In comparison, the baseline accuracy for the majority class (that is, the most frequent label among the three categories (A–C) within each validation set) is 87% for post titles, 49% for comments made by the post’s author, and 55% for comments from other accounts. Note that the majority class differs across the three validation sets: for post titles and comments made by the post’s author, the majority label is Implies AI involvement, while for comments from other accounts, the majority label is Does not imply AI.
@@
Following validation, we classified all data entities in the dataset using these prompts, following a process inspired by the QuaLLM framework (Rao et al., 2024). According to this process we submitted the prompt, consisting in each case of the relevant class explanation and the in-context examples, with every 100 input texts at a time [^6] Note that during the validation phase we also compared the accuracy values obtained by varying the batch size (i.e., 10 instead of 100 texts at a time) and achieved identical results.. Similarly, we requested that the LLM return the final output as a dictionary, including both the reasoning and classifications for all these 100 texts.
@@ 6.2. Participation and Newcomers
The previous sections focused on the volume of posts and discussions over time and in relation to other image posts in the same communities. In this section, we address our second research question by examining participation patterns in the creation of AI-related posts across the different art communities. In particular, we focus on newcomers – users who make their first-ever image-sharing post to a certain subreddit. Is there evidence that AI-generated images make it easier for newcomers to contribute content – and if so, do these newcomers stay and contribute over time? Does this pattern diverge for those who are making Transparent AI posts to AI-disallowed versus AI-neutral communities?
Overall, the participation of newcomers was significantly more emphasized in making Transparent AI posts than the overall presence of newcomers in posting images to these communities. In AI-disallowed communities, newcomers usually publish 6.71% of all image-based posts. This ratio is similar in AI-neutral communities, with 6.51% of newcomers publishing image-based posts. However, newcomers made as much as 25.15% of Transparent AI posts in AI-disallowed communities during our period of analysis. In AI-neutral communities, newcomers made a similar 22.99% of the Transparent AI posts. Chi-square tests showed that the proportion of posts made by newcomers is significantly different between Transparent AI posts and image-based posts (AI-disallowed communities: $\chi^{2}(1,N=3,964,964)=1796.71,p<0.00001$; AI-neutral communities: $\chi^{2}(1,N=2,328,041)=521.10,p<0.00001$). This over-representation in newcomers posting AI-generated content suggests that this kind of content may encourage and broaden community participation.
@@
Overall, the participation trends show that newcomers more heavily rely on posting AI content to participate in these communities, but that (as expected) their participation does not last long in AI-disallowed communities (see section 7 for additional discussion). These trends may relate to the type of interactions around AI-generated content, which we look at next.
@@ 6.3. Feedback, Interactions and Sentiment
The time dynamics show a different response to suspected images in the two types of communities. Furthermore, to better understand the response, we conducted a sentiment analysis of comments suggesting AI involvement in the Suspected AI posts. To label the sentiment of comments implying AI, we used a large language model as a classifier, following a similar approach to that described in the previous sections, which is based on prior studies (Wei et al., 2022; Rao et al., 2024). We used the LLM classifier to label the comments as Negative, Neutral, or Positive, using a prompt that can be found in Appendix D, along with examples of input texts and their corresponding classifications. To assess the accuracy of this method, we manually labeled a random sample of 100 comments and achieved an accuracy rate of 92% for this sentiment classification task.
Using these labels, we looked at differences in comment sentiment between AI-disallow and AI-neutral communities, and how they may have changed over time. To do that, we compared sentiment in two distinct time periods: January-March 2023 and October-December 2023 (the end of our data). In January-March 2023, the percentage of negative sentiment in comments in AI-disallowing communities was 39% (6% were positive, and the rest neutral). The ratio of negative sentiment was lower in AI-neutral communities, at around 28%. By October-December 2023, however, the pattern reversed, with AI-neutral communities exhibiting a higher percentage of negative comments (51%) than AI-disallowing communities (42%). The reason for the flip could be the better control exercised by the AI-disallowing communities about AI posts, or a normative shift in the AI-neutral communities with respect to AIGC. Still, the data shows an increase over time in the proportion of comments expressing negative sentiment among those claiming AI involvement, across both community categories.
@@ 7. Discussion
The data processing and analysis steps may also have certain limitations, which could be addressed in future research. Identifying posts containing AI-generated media relies on text-based methods, which sometimes may capture user opinion and reactions, rather than accurately reflect the actual prevalence of AI-generated content. For example, in AI-neutral subreddits where AI-generated content is permitted, users can be less inclined to explicitly mention it in posts or comments. In contrast, in AI-disallowed subreddits, users may be more likely to claim that others’ content is AI-generated as it violates community rules. In the categorization of the communities step, given that AI-neutral communities do not explicitly address AI usage in their guidelines or descriptions, while AI-disallowed communities explicitly prohibit it, the observed differences in trends between these groups may reflect broader factors, such as more adaptive approach to emerging technologies by the AI-disallowed communities. That is, factors beyond the subreddit-specific policies regarding AI-generated content. Additionally, the analysis may face potential limitations related to the identification of newcomers, as the dataset may include bot or troll accounts considered as newcomers. Bots, in particular, are more likely to use AI-generated content (Grimme et al., 2022; Ferrara, 2023), and may be at higher risk of suspension by the platform. During data collection, measures were taken to filter out potential bot accounts based on usernames and repeated comments. However, these efforts may not have entirely excluded all accounts exhibiting bot-like behavior. Finally, future work could refine the sentiment analysis by moving beyond the three-label scheme (Negative, Neutral, Positive) to provide more nuanced insights.